OpenStudy (anonymous):

The vertex of the parabola, given by the equation h (x) = cx^2 − 2x + 1, is

OpenStudy (across):

\[h(x)=cx^2-2x+1,\]\[h'(x)=2cx-2=0\implies x=\frac{1}{c},\]\[h(x)=-1/c+1,\]\[\left ( \frac{1}{c},1-\frac{1}{c} \right ).\]